# Exact Diagonalization, Matrix Product States and Conformal Perturbation Theory Study of a 3D Ising Fuzzy Sphere Model

### v1

@@ 3 Conformal Perturbation Theory
We will next review the effective field theory (EFT) which will allow us to quantify the deviation between microscopic and exact CFT energy levels, i.e. the error in (2) . We assume that the microscopic theory has a local Hamiltonian, as is true for 1+1D spin chains and for 2+1D fuzzy sphere regulator models. The basic RG intuition suggests that any local theory close to a CFT can be described by an EFT perturbing the CFT by a collection of local operators Zamolodchikov:1986gt ; Zamolodchikov:1987ti .
@@ 3.1 Corrections to correlation functions in $\mathbb{R}^{d}$
where ${\cal V}$ may be any CFT operator allowed by the microscopic symmetry (excluding total derivatives as having no effect on a manifold without a boundary), and $G_{\cal V}$ are the couplings. They depend on the microscopic theory and in practice their values are unknown. We assume that the perturbation arises at the microscopic distance scale $a=1$, a short distance cutoff which sets our unit of length. Conformal perturbation theory Zamolodchikov:1986gt ; Zamolodchikov:1987ti ; Cappelli:1989yu ; Cappelli:1991ke ; Guida:1995kc ; Cardy:1996xt is a well-known method for evaluating correlators starting from (13). The idea is to insert $e^{-\Delta S}$ inside the CFT correlators, expand the exponential, and evaluate the resulting integrals in $x$. For example, the two point function of a field $\Phi$ would be evaluated in this approach as:
@@
$G_{\cal V}\xi^{d-\Delta_{\cal V}}\sim 1\qquad\text{(${\cal V}$ relevant)}\,.$) . (15)
@@ 3.2 Corrections to the spectrum on the sphere
To evaluate the energy levels, it’s instructive to pass to the Hamiltonian formalism. The Hamiltonian corresponding to the action (16) is obtained as the sum of the CFT Hamiltonian on $S^{d-1}_{R}$ and of the perturbing term, which is the $\tau$-integrand in (16) evaluated at $\tau=0$ Yurov:1989yu :
@@
The perturbation is assumed to be spherically symmetric, as is appropriate for the description of the fuzzy sphere model which has exact spherical symmetry at the microscopic level. [^10] If instead only a discrete subgroup $\Gamma$ of SO$(d)$ is preserved, the couplings $G_{\cal O}$ may have nontrivial dependence on the sphere coordinates, consistent with $\Gamma$-invariance, as in Lao:2023zis where $\Gamma$ was the icosahedral subgroup of SO(3).
@@
We could obtain nonperturbative results for the energy levels by diagonalizing the infinite matrix $H_{\rm CFT}+V$, e.g. truncating to finite size and extrapolating, referred to as the Truncated Conformal Space Approach Yurov:1989yu ; Hogervorst:2014rta ; James:2017cpc . This would be needed for $R\gtrsim\xi$, when corrections to CFT energy levels are significant. However, since in this paper we work at $1\ll R\ll\xi$, perturbation theory will be sufficient for our purposes. Indeed, in this case the couplings $g_{\cal V}$ in (19) are all small:
@@
For higher-order corrections, it would be actually advantageous to go back to the Lagrangian description, reducing them to integrals of correlation functions with multiple operator insertions on the cylinder Ludwig:1987gs ; Reinicke1987a . [^11] Such expressions compactly encode the infinite sums over intermediate states which one would obtain if one were to use higher-order Schroedinger-Rayleigh perturbation theory, also taking into account constraints from conformal symmetry on the matrix elements. Depending on the dimension of the perturbing operator, this may give UV divergences, regulated by the short-distance cutoff $a$, and requiring renormalization of the couplings EliasMiro:2021aof ; EliasMiro2023 . In this paper we will be mostly content with the first-order correction (21) for which these issues do not arise.
@@ 3.3 Matrix elements from the OPE coefficients
We now discuss how the matrix elements (21) are computed (see e.g. Lao:2023zis ). The basic idea is to represent the matrix element as the (integral of) correlator on the cylinder where the states $|\Phi\rangle$ and $\langle\Phi|$ are prepared by inserting the operators $\Phi_{\rm cyl}$ in infinite past and $(\Phi_{\rm cyl})^{\dagger}$ in infinite future. Then, one performs the Weyl transformation from the cylinder back to $\mathbb{R}^{d}$, which maps the cylinder correlator to a CFT three-point function in $\mathbb{R}^{d}$ with operators $\Phi$ and $\Phi^{\dagger}$ inserted at 0 and $\infty$. We will consider several cases of increasing degree of complexity.
@@
Here we used Polyakov’s formula for the scalar primary three-point function Polyakov:1970xd , which implies $\langle{\cal O}(\infty){\cal V}(n){\cal O}(0)\rangle=f_{{\cal O}{\cal O}{\cal V}},$ where $f_{{\cal O}{\cal O}{\cal V}}$ is the OPE coefficient in the standard normalization. [^12] OPE coefficients of scalar primaries are normalized via ${\cal O}_{1}(x){\cal O}_{2}(0)\supset f_{123}|x|^{\Delta_{3}-\Delta_{1}-\Delta_{2}}{\cal O}_{3}(0)$. This is the normalization used in all the literature. When one or more of the operators carry Lorentz indices, there is no commonly accepted convention. Our convention will be explained when needed. Note that the result is spherically symmetric as expected. Note also that the state is unit normalized: $\langle{\cal O}|{\cal O}\rangle=\langle{\cal O}(\infty){\cal O}(0)\rangle=1$ working in the usual normalization of scalar primaries. The final result for the energy correction is given by (the sum over ${\cal V}$ is implied)
@@
The descendants of ${\cal O}$ at level $k$ are degenerate of unperturbed energy $\Delta_{\cal O}+k$. They form a symmetric tensor product representation ${\rm Sym}^{k}(\mathbf{d})$ where $\mathbf{d}$ is the fundamental of SO$(d)$. This representation is reducible for $k\geqslant 2$, and its irreducible components have different relative factors, so that the degeneracy is lifted by the perturbation. We describe here the result for $k=1$, $d=3$. The $k=1$ descendants form a vector representation $\mathbf{3}$ of SO$(3)$ which is irreducible, and remains degenerate under the spherically symmetric perturbation. We may use (21) but we have to remember to normalize the states. The matrix element and the norm come out to be:
@@
where the notation $A(k,\rho)$ means that we are dealing with the correction for level $k$ descendants transforming in the irrep $\rho$ of SO(3). [^13] Irreps of SO(3) will be denoted interchangeably by spin in parentheses, or multiplicity in boldface. E.g. spin-1 irrep will be denoted as $(\ell=1)$ or $\mathbf{3}$. Such relations between corrections for primaries and descendants mean that when CFT is perturbed, many energy levels will move in a correlated fashion. The derivation of (26) and analogous relations for descendants up to $k=4$ are discussed in App. B and in the notebook notebook .
@@
a $-\ell$ primary operator ${\cal U}^{(\ell)}$ contracted with a coefficient tensor $t$ to get a rotationally invariant object. [^14] As noted after (13) we do not have to consider total derivatives. This is obvious from the Lagrangian point of view. In the Hamiltonian formalism, total derivative perturbations have vanishing matrix elements at first order. In higher orders they are redundant—can be removed by a similarity transformation of the Hamiltonian leaving the spectrum invariant (Lao:2023zis, , footnote 7). See (Reinicke1987b, , Eq. (3.9)) for an example of the latter. The coefficient tensor $t$ in (27) must be built out of $\delta_{\mu\nu}$ and of the unit vector $n_{\mu}\in S^{d-1}$. The primaries being symmetric traceless tensors, the only nonvanishing contraction is
@@
where $(x)_{\ell}$ is the Pochhammer symbol. To derive this equation one needs to use the form of the three-point function scalar-scalar-(spin $\ell$), see e.g. (Poland:2018epd, , Eq. (23)). The correction to the energy of the primary state is given by an equation analogous to (23):
@@
A comment is in order concerning normalization of the OPE coefficients $f_{{\cal O}{\cal O}{\cal U}^{(\ell)}}$. In this paper we use the normalization as in Simmons-Duffin2017 , which corresponds to the OPE ${\cal O}_{1}(x){\cal O}_{2}(0)\supset 2^{\ell/2}f_{{\cal O}_{1}{\cal O}_{2}{\cal U}^{(\ell)}}|x|^{\Delta_{\cal U}-\Delta_{{\cal O}_{1}}-\Delta_{{\cal O}_{2}}}x_{\mu_{1}}\cdots x_{\mu_{\ell}}{\cal U}^{(\ell)}_{\mu_{1}\ldots\mu_{\ell}}(0)$. In this normalization the three-point function is $2^{\ell/2}$ times the expression given in (Poland:2018epd, , Eq. (23)), which does not contain $\ell$-dependent factors. This explains the factor $2^{\ell/2}$ in (29). The rest of $q_{\ell}$ in (29) comes from subtracting traces in the three-point function.
@@ 4 Warmup: CFT and CPT for a 1+1D Ising spin chain
In order to illustrate the power of CPT in combination with numerical energy spectra, we study a well known 1+1D lattice model in this section—the antiferromagnetic Ising model in a transverse and a longitudinal magnetic field—which features a line of Ising critical points in the magnetic fields plane. CPT results for the perturbed 2D Ising CFT were derived in two papers by Reinicke Reinicke1987a ; Reinicke1987b almost fourty years ago. Together these data and results allow to infer the couplings constants perturbing the CFT at and in the vicinity of the critical point for finite size systems. Furthermore we can analyze the finite-size scaling of the inferred couplings constants and we can confront them to their expected RG scaling behaviour.
@@
i.e. the antiferromagnetic Ising model in a uniform longitudinal ($h_{z}$) and transverse field ($h_{z}$). At $h_{z}=0$ this model is unitarily equivalent to the standard ferromagnetic transverse field Ising model, which can be solved exactly using a mapping to fermions Pfeuty1970 , with a critical point at $h_{x}=1$. For $h_{z}\in(-2,2)$ the model has a line of critical Ising points at values of $h_{x}$ which depend on $h_{z}$ and are not known exactly, as the model loses its integrability away from $h_{z}=0$. For illustration purposes we will focus on the value $h_{z}=1$ only and explore the critical point and its vicinity by tuning $h_{x}$. The model has lattice translation symmetry and spatial reflection symmetry with respect to a site (or the center of a bond). For $h_{z}\neq 0$, there is no microscopic on-site $\mathbb{Z}_{2}$ symmetry. Instead, the global $\mathbb{Z}_{2}$ symmetry of the Ising CFT describing the critical point corresponds in the microscopic model to the translation by one lattice spacing. (It is the same $\mathbb{Z}_{2}$ which breaks spontaneously in the Néel phase of our model.)
Figure 5: Comparison of the 2D Ising CFT spectrum with the ED energy gaps. Circles and full lines: ED energy gaps for the Hamiltonian (32) for $N=26$ rescaled by $R/v$ and plotted as a function of $h_{x}$ for $h_{z}=1$. Squares and dashed lines: Reinicke CPT model obtained by fit to $T,\sigma,\partial\sigma,\varepsilon,\partial\varepsilon$. The vertical dashed line denotes the location of the critical point. The CPT model describes the ED energy gaps very well, despite the visible deviations of the ED energy gaps from the 2D Ising CFT values.
@@
In what follows we will show that a much improved agreement between the ED and the CFT can be obtained by fitting the rescaled ED energy gaps to CPT predictions for the energy levels of 2D Ising CFT perturbed by relevant and irrelevant perturbations, obtained by Reinicke Reinicke1987a ; Reinicke1987b and reviewed in App. C.
As a first step we track the lowest two $\mathbb{Z}_{2}$-odd energy levels at momentum 00 and $2\pi/R$, [^16] On the lattice these two levels have a momentum offset of $\pi/a$ because the magnetic order is antiferromagnetic in the model considered here. The $\mathbb{Z}_{2}$ quantum number is therefore encoded in whether the energy levels are found close to lattice momentum 00 ($\mathbb{Z}_{2}$-even) or $\pi/a$ ($\mathbb{Z}_{2}$-odd). which correspond to $\sigma$ and $\partial\sigma$. We then use the Reinicke formula to first order in $g_{\varepsilon}$ to infer the speed of light $v$ and the value of $g_{\varepsilon}$ for all values of $h_{x}$ shown.
@@
Figure 7: Analysis of the ED energy gaps with the help of a more sophisticated conformal perturbation theory model perturbing the 2D Ising CFT by $\varepsilon,T\bar{T},T^{2}+\bar{T}^{2}$. Top row: extracted $g_{\varepsilon}$ coupling from the full CPT model (left) and the residual differences between the CFT model at that order and the considered ED levels (right). Middle row: extracted couplings $g_{T\bar{T}}$ (left) and $g_{T^{2}+\bar{T}^{2}}$ (right). Bottom row: the extracted couplings $g_{T\bar{T}}$ and $g_{T^{2}+\bar{T}^{2}}$ (left) and $g_{\epsilon}$ (right) at $h$ near $h_{c}$, plotted as a function of system size $N$. The dotted lines illustrate the scaling expected from the RG.
@@
In the last step we include in the analysis the CPT predictions for the perturbations by two leading irrelevant parity-invariant operators of scaling dimensions 4: $T\bar{T}$ and $T^{2}+\bar{T}^{2}$. We do so by fitting the CPT expressions for the ED energy levels $T,\sigma,\partial\sigma,\varepsilon,\partial\varepsilon$ with $v$, $g_{\varepsilon}$, $g_{T\bar{T}}$ and $g_{T^{2}+\bar{T}^{2}}$ as parameters. The results are shown in the first two rows of Fig. 7. While $g_{\varepsilon}(h_{x})$ is almost unchanged compared to Fig. 6 on the scale of the plot, it is notable that the residual difference between the CPT and the ED levels dropped by more than a factor 20 by including the irrelevant operators in the fit. The extracted new coupling constants are shown in the middle row. For our chosen microscopic model the two couplings are both of negative sign, and of comparable magnitude. [^17] It is known that for our model at $h_{z}=0$ the extracted coupling $g_{T\bar{T}}$ would be zero to the considered order, while $g_{T^{2}+\bar{T}^{2}}$ is also negative Reinicke1987b . In the free fermion case $h_{z}=0$ this is due to lattice induced curvature of the dispersion deviating from linearity. While $T\bar{T}$ is the known leading scalar primary, which controls correction to scaling in the 2D Ising universality class, the presence of the spin-4 operator ${T^{2}+\bar{T}^{2}}$ is due to presence of Lorentz symmetry breaking terms (e.g. a square lattice in 2D or a Hamiltonian formulation in 1+1D, leading to a space-time anisotropy). Furthermore the negative value of its coupling in our model is likely a band curvature effect which can be understood in the limit $h_{z}=0$. Analogues of the two irrelevant operators will also play a role on our analysis of the fuzzy sphere geometry. In the bottom row we finally show the finite size behavior of the two extracted couplings $g_{T\bar{T}}$ and $g_{T^{2}+\bar{T}^{2}}$ at the critical point. Since these operators both have scaling dimensions 4, we would expect them to flow away with an exponent $2-4=-2$ with increasing system size $N$, and the data shown in the bottom row nicely agrees with this expectation. Similarly, the value of $g_{\varepsilon}$ in Figs. 6(a) and 7(a) can be seen to increase approximately linearly with the system size away from criticality, in agreement with $2-1=1$.

### v2

@@ 3 Conformal Perturbation Theory
We will next review the effective field theory (EFT) which will allow us to quantify the deviation between microscopic and exact CFT energy levels, i.e. the error in (2) . We assume that the microscopic theory has a local Hamiltonian, as is true for 1+1D spin chains and for 2+1D fuzzy sphere regulator models. The basic RG intuition suggests that any local theory close to a CFT can be described by an EFT perturbing the CFT by a collection of local operators [50, 51].
@@ 3.1 Corrections to correlation functions in $\mathbb{R}^{d}$
where $\cal V$} may be any CFT operator allowed by the microscopic symmetry (excluding total derivatives as having no effect on a manifold without a boundary), and $G_{\cal V}$ are the couplings. They depend on the microscopic theory and in practice their values are unknown. We assume that the perturbation arises at the microscopic distance scale $a=1$, a short distance cutoff which sets our unit of length. Conformal perturbation theory [50, 51, 52, 53, 54, 55] is a well-known method for evaluating correlators starting from (13). The idea is to insert $e^{-\Delta S}$ inside the CFT correlators, expand the exponential, and evaluate the resulting integrals in $x$. For example, the two point function of a field $\Phi$ would be evaluated in this approach as:
@@
$G_{\cal V}\xi^{d-\Delta_{\cal V}}\sim 1\qquad\text{(${\cal V}$ relevant)}\,.$ (15)
@@ 3.2 Corrections to the spectrum on the sphere
To evaluate the energy levels, it’s instructive to pass to the Hamiltonian formalism. The Hamiltonian corresponding to the action (16) is obtained as the sum of the CFT Hamiltonian on $S^{d-1}_{R}$ and of the perturbing term, which is the $\tau$-integrand in (16) evaluated at $\tau=0$ [56]:
@@
Here $\mathcal{V}_{\rm cyl}$ is the same CFT operator as in (16). [^10] From perturbative field theory point of view, it may look like operators containing time derivatives change when going to the Hamiltonian formalism, due to the Legendre transform. But this change only affects the operator’s expression in terms of fundamental fields. The matrix elements and correlators are unchanged, so from the CFT point of view, it is the same operator and is denoted accordingly. The perturbation is assumed to be spherically symmetric, as is appropriate for the description of the fuzzy sphere model which has exact spherical symmetry at the microscopic level. [^11] If instead only a discrete subgroup $\Gamma$ of SO(d)(d) is preserved, the couplings $G_{\cal O}$ may have nontrivial dependence on the sphere coordinates, consistent with $\Gamma$-invariance, as in [29] where $\Gamma$ was the icosahedral subgroup of SO(3).
@@
[…]
@@
For higher-order corrections, it would be actually advantageous to go back to the Lagrangian description, reducing them to integrals of correlation functions with multiple operator insertions on the cylinder [59, 32]. [^12] Such expressions compactly encode the infinite sums over intermediate states which one would obtain if one were to use higher-order Schroedinger-Rayleigh perturbation theory, also taking into account constraints from conformal symmetry on the matrix elements. Depending on the dimension of the perturbing operator, this may give UV divergences, regulated by the short-distance cutoff $a$, and requiring renormalization of the couplings [60, 61]. In this paper we will be mostly content with the first-order correction (21) for which these issues do not arise.
@@ 3.3 Matrix elements from the OPE coefficients
We now discuss how the matrix elements (21) are computed (see e.g. [29]). The basic idea is to represent the matrix element as the (integral of) correlator on the cylinder where the states $|\Phi\rangle$ and $\langle\Phi|$ are prepared by inserting the operators $\Phi_{\rm cyl}$ in infinite past and $(\Phi_{\rm cyl})^{\dagger}$ in infinite future. Then, one performs the Weyl transformation from the cylinder back to $\mathbb{R}^{d}$, which maps the cylinder correlator to a CFT three-point function in $\mathbb{R}^{d}$ with operators $\Phi$ and $\Phi^{\dagger}$ inserted at 0 and $\infty$. We will consider several cases of increasing degree of complexity.
@@
Here we used Polyakov’s formula for the scalar primary three-point function [62], which implies $\langle{\cal O}(\infty){\cal V}(n){\cal O}(0)\rangle=f_{{\cal O}{\cal O}{\cal V}},$ where $f_{{\cal O}{\cal O}{\cal V}}$ is the OPE coefficient in the standard normalization. [^13] OPE coefficients of scalar primaries are normalized via ${\cal O}_{1}(x){\cal O}_{2}(0)\supset f_{123}|x|^{\Delta_{3}-\Delta_{1}-\Delta_{2}}{\cal O}_{3}(0)$. This is the normalization used in all the literature. When one or more of the operators carry Lorentz indices, there is no commonly accepted convention. Our convention will be explained when needed. Note that the result is spherically symmetric as expected. Note also that the state is unit normalized: $\langle{\cal O}|{\cal O}\rangle=\langle{\cal O}(\infty){\cal O}(0)\rangle=1$ working in the usual normalization of scalar primaries. The final result for the energy correction is given by (the sum over $\cal V$} is implied)
@@
[…]
@@
where the notation $A(k,\rho)$ means that we are dealing with the correction for level $k$ descendants transforming in the irrep $\rho$ of SO(3). [^14] Irreps of SO(3) will be denoted interchangeably by spin in parentheses, or multiplicity in boldface. E.g. spin-1 irrep will be denoted as $(\ell=1)$ or $\mathbf{3}$. Such relations between corrections for primaries and descendants mean that when CFT is perturbed, many energy levels will move in a correlated fashion. The derivation of (26) and analogous relations for descendants up to $k=4$ are discussed in App. B and in the notebook [34].
@@
a spin-$\ell$ primary operator ${\cal U}^{(\ell)}$ contracted with a coefficient tensor $t$ to get a rotationally invariant object. [^15] As noted after (13) we do not have to consider total derivatives. This is obvious from the Lagrangian point of view. In the Hamiltonian formalism, total derivative perturbations have vanishing matrix elements at first order. In higher orders they are redundant—can be removed by a similarity transformation of the Hamiltonian leaving the spectrum invariant [29, footnote 7]. See [33, Eq. (3.9)] for an example of the latter. The coefficient tensor $t$ in (27) must be built out of $\delta_{\mu\nu}$ and of the unit vector $n_{\mu}\in S^{d-1}$. The primaries being symmetric traceless tensors, the only nonvanishing contraction is
@@
where $(x)_{\ell}$ is the Pochhammer symbol. To derive this equation one needs to use the form of the three-point function scalar-scalar-(spin $\ell$), see e.g. [37, Eq. (23)]. The correction to the energy of the primary state is given by an equation analogous to (23):
@@
A comment is in order concerning normalization of the OPE coefficients $f_{{\cal O}{\cal O}{\cal U}^{(\ell)}}$. In this paper we use the normalization as in [13], which corresponds to the OPE ${\cal O}_{1}(x){\cal O}_{2}(0)\supset$ $2^{\ell/2}f_{{\cal O}_{1}{\cal O}_{2}{\cal U}^{(\ell)}}|x|^{\Delta_{\cal U}-\Delta_{{\cal O}_{1}}-\Delta_{{\cal O}_{2}}}x_{\mu_{1}}\cdots x_{\mu_{\ell}}{\cal U}^{(\ell)}_{\mu_{1}\ldots\mu_{\ell}}(0)$. In this normalization the three-point function is $2^{\ell/2}$ times the expression given in [37, Eq. (23)], which does not contain $\ell$-dependent factors. This explains the factor $2^{\ell/2}$ in (29). The rest of $q_{\ell}$ in (29) comes from subtracting traces in the three-point function.
@@ 4 Warmup: CFT and CPT for a 1+1D Ising spin chain
In order to illustrate the power of CPT in combination with numerical energy spectra, we study a well known 1+1D lattice model in this section—the antiferromagnetic Ising model in a transverse and a longitudinal magnetic field—which features a line of Ising critical points in the magnetic fields plane. CPT results for the perturbed 2D Ising CFT were derived in two papers by Reinicke [32, 33] almost fourty years ago. Together these data and results allow to infer the couplings constants perturbing the CFT at and in the vicinity of the critical point for finite size systems. Furthermore we can analyze the finite-size scaling of the inferred couplings constants and we can confront them to their expected RG scaling behaviour.
@@
i.e. the antiferromagnetic (J>0) Ising model in a uniform longitudinal ($h_{z}$) and transverse field ($h_{z}$). At $h_{z}=0$ this model is unitarily equivalent to the standard ferromagnetic transverse field Ising model, which can be solved exactly using a mapping to fermions [63], with a critical point at $h_{x}=1$. For $h_{z}\in(-2,2)$ the model has a line of critical Ising points at values of $h_{x}$ which depend on $h_{z}$ and are not known exactly, as the model loses its integrability away from $h_{z}=0$. For illustration purposes we will focus on the value $h_{z}=1$ only and explore the critical point and its vicinity by tuning $h_{x}$. The model has lattice translation symmetry and spatial reflection symmetry with respect to a site (or the center of a bond). For $h_{z}\neq 0$, there is no microscopic on-site $\mathbb{Z}_{2}$ symmetry. Instead, the global $\mathbb{Z}_{2}$ symmetry of the Ising CFT describing the critical point corresponds in the microscopic model to the translation by one lattice spacing. (It is the same $\mathbb{Z}_{2}$ which breaks spontaneously in the Néel phase of our model.)
Figure 5: Comparison of the ED energy gaps with the 2D Ising CFT spectrum and with the 2D Ising spectrum corrected with the help of a CPT model. Circles and full colored lines: ED energy gaps for the Hamiltonian (32) for $N=26$ rescaled by $R/v$ and plotted as a function of $h_{x}$ for $h_{z}=1$. Horizontal gray lines: the exact 2D Ising CFT spectrum. Squares and dashed colored lines: the 2D Ising spectrum corrected by the Reinicke CPT model with couplings $g_{\varepsilon}$, $g_{T\widetilde{T}}$, $g_{T^{2}+\widetilde{T}^{2}}$ obtained by fit to $T,\sigma,\partial\sigma,\varepsilon,\partial\varepsilon$ levels. The vertical dashed line denotes the location of the critical point, determined from the equation $g_{\varepsilon}=0$. The CPT model describes the ED energy gaps very well, despite the visible deviations of the ED energy gaps from the 2D Ising CFT values. The CPT and ED agreement is so good that the dashed lines are essentially invisible, hidden behind solid lines, for all the levels except $\partial^{2}\varepsilon$. Accordingly the circles are always inside the squares.
@@
In what follows we will show that a much improved agreement between the ED and the CFT can be obtained by fitting the rescaled ED energy gaps to CPT predictions for the energy levels of 2D Ising CFT perturbed by relevant and irrelevant perturbations, obtained by Reinicke [32, 33] and reviewed in App. C.
As a first step we track the lowest two $\mathbb{Z}_{2}$-odd energy levels at momentum 0 and $2\pi/N=1/R$, [^17] On the lattice these two levels have a momentum offset of $\pi/a$ because the magnetic order is antiferromagnetic in the model considered here. The $\mathbb{Z}_{2}$ quantum number is therefore encoded in whether the energy levels are found close to lattice momentum 0 ($\mathbb{Z}_{2}$-even) or $\pi/a$ ($\mathbb{Z}_{2}$-odd). Note that we consider even-length chains for simplicity. which correspond to $\sigma$ and $\partial\sigma$. We then use the Reinicke formula to first order in $g_{\varepsilon}$ to infer the speed of light $v$ and the value of $g_{\varepsilon}$ for all values of $h_{x}$ shown.
@@
Figure 7: Analysis of the ED energy gaps with the help of a more sophisticated conformal perturbation theory model perturbing the 2D Ising CFT by $\varepsilon,T\bar{T},T^{2}+\bar{T}^{2}$. Top row: extracted $g_{\varepsilon}$ coupling from the full CPT model (left) and the residual differences between the CFT model at that order and the considered ED levels (right). Middle row: extracted couplings $g_{T\bar{T}}$ (left) and $g_{T^{2}+\bar{T}^{2}}$ (right). Bottom row: the extracted couplings $g_{T\bar{T}}$ and $g_{T^{2}+\bar{T}^{2}}$ (left) and $g_{\epsilon}$ (right) at $h$ near $h_{c}$, plotted as a function of system size NN. The dotted lines illustrate the scaling expected from the RG.
@@
In the last step we include in the analysis the CPT predictions for the perturbations by two leading irrelevant parity-invariant operators of scaling dimensions 4: $T\bar{T}$ and $T^{2}+\bar{T}^{2}$. We do so by fitting the CPT expressions for the ED energy levels $T,\sigma,\partial\sigma,\varepsilon,\partial\varepsilon$ with $v$, $g_{\varepsilon}$, $g_{T\bar{T}}$ and $g_{T^{2}+\bar{T}^{2}}$ as parameters. The results are shown in the first two rows of Fig. 7. While $g_{\varepsilon}(h_{x})$ is almost unchanged compared to Fig. 6 on the scale of the plot, it is notable that the residual difference between the CPT and the ED levels dropped by more than a factor 20 by including the irrelevant operators in the fit. The extracted new coupling constants are shown in the middle row. For our chosen microscopic model the two couplings are both of negative sign, and of comparable magnitude. [^18] It is known that for our model at $h_{z}=0$ the extracted coupling $g_{T\bar{T}}$ would be zero to the considered order, while $g_{T^{2}+\bar{T}^{2}}$ is also negative [33]. In the free fermion case $h_{z}=0$ this is due to lattice induced curvature of the dispersion deviating from linearity. While $T\bar{T}$ is the known leading scalar primary, which controls correction to scaling in the 2D Ising universality class, the presence of the spin-4 operator $T^{2}+\bar{T}^{2}$} is due to presence of Lorentz symmetry breaking terms (e.g. a square lattice in 2D or a Hamiltonian formulation in 1+1D, leading to a space-time anisotropy). Furthermore the negative value of its coupling in our model is likely a band curvature effect which can be understood in the limit $h_{z}=0$. Analogues of the two irrelevant operators will also play a role on our analysis of the fuzzy sphere geometry. In the bottom row we finally show the finite size behavior of the two extracted couplings $g_{T\bar{T}}$ and $g_{T^{2}+\bar{T}^{2}}$ at the critical point. Since these operators both have scaling dimensions 4, we would expect them to flow away with an exponent 24=22-4=-2 with increasing system size $N$, and the data shown in the bottom row nicely agrees with this expectation. Similarly, the value of $g_{\varepsilon}$ in Figs. 6(a) and 7(a) can be seen to increase approximately linearly with the system size away from criticality, in agreement with 21=12-1=1.
As one can see in Fig. 5, this last CPT model describes the ED spectrum extremely well, so well that deviation is invisible by naked eye for all levels except for $\partial^{2}\varepsilon$.